OpenStudy (anonymous):

find indefinite integral (5 dx)/((x)ln(5 x))

OpenStudy (anonymous):

u sub for this one

OpenStudy (anonymous):

\[5\int\frac{dx}{\ln(5x)}\] \[u=\ln(x), du =\frac{1}{x}dx\] \[5\int\frac{du}{u}=5\ln(u) \] etc

OpenStudy (dumbcow):

looks like it will turn out to be a nested log

OpenStudy (anonymous):

oops second line should be \[u=\ln(5x), du =\frac{1}{x}dx\] but it makes no difference

OpenStudy (anonymous):

"nested log" maybe owlfred can make a home there

OpenStudy (dumbcow):

:)

OpenStudy (anonymous):

(5 dx)/((x)ln(5 x)) Let u = ln 5x => du = 5dx/ 5x = dx/x

OpenStudy (anonymous):

-> 5 integral du / u = 5 lnu +c = 5 ln [ ln(5x) ] + c